OpenStudy (lisa123):

Determine whether the vectors u and v are parallel, orthogonal, or neither u= <6,-2> v=<2,6>

OpenStudy (shamim):

If u and v are parallel then the cross product of u and v will b =0

OpenStudy (lisa123):

I think their orthogonal

OpenStudy (shamim):

u cross v=?

OpenStudy (shamim):

Orthogonal=perpendicular???

OpenStudy (shamim):

Ya they r orthogonal

OpenStudy (lisa123):

6(2) + -2(6) 12 + -12= 0

OpenStudy (shamim):

Bt hw do u know they r orthogonal?

OpenStudy (shamim):

Ivthink u calculated dot product

OpenStudy (shamim):

U got u dot v=0

OpenStudy (shamim):

Right?

OpenStudy (lisa123):

I don't understand

OpenStudy (shamim):

Ya

OpenStudy (lisa123):

Really? Are u sure?

OpenStudy (shamim):

Openstudy website disturbing little!!

OpenStudy (shamim):

My answers r not in orders!!

OpenStudy (lisa123):

I thought orthogonal meant when the final result =0

OpenStudy (shamim):

Latest answers going to the top!

OpenStudy (shamim):

Anyway enjoying helping u!!!

OpenStudy (shamim):

Actually angle between ur given 2 vectors will b 90 degree. Its called orthogonal

OpenStudy (shamim):

C plz

OpenStudy (shamim):

u dot v=0 u*v*cos theta=0 Cos theta=0/uv=0 Theta=90degree

OpenStudy (shamim):

Glad to help u!!!